simplify

x^2 - 8x plus 7x - 42
---------------divide----------
x^2 + 12x + 36 plus x^2 +12x +36

Answer by london maths tutor(243)

(x^2 - 8x + 7x - 42)/(x^2+12x+36+x^2+12x+36)
=(x^2-x-42)/[2*(x^2+12x+36)]
Factorise the various bracket:
x^2-x-42 = (x-6)(x+7)
(x^2+12x+36) = (x+6)(x+6)
Put the factorised brackets into the question:
[(x-6)(x-7)]/[2*(x+6)(x+6)]